find the sum to n terms of 1+3+32+.......+3n-1?
Pls help me with detailed solution

Grade:11

1 Answers

Arun
5 years ago
Dear Surya
 
If we look carefully, it is a G.P. of n terms with first term as 3^0 = 1
common ratio = 3
 
Now
 
sum = 1 * (3^n   – 1)/(3 – 1)
 
 = (3^n  – 1 )/ 2